# EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN) CERN-EP-2023-280 LHCb-PAPER-2023-037

### v1

@@
Submitted to Eur. Phys. J. C
© 2023 CERN for the benefit of the LHCb collaboration. CC BY 4.0 licence.
@@ 1 Introduction
where indicates a yield, indicates the efficiency determined from simulation with data-driven corrections, and is the known branching fraction of the decay [21].
@@ 2 Detector and simulation
Simulation is used to optimise the event selection procedure, to model the shape of the candidate invariant-mass distributions and to estimate efficiencies accounting for the effects of the detector acceptance, reconstruction and selection criteria. In the simulation, collisions are generated using Pythia [31] with a specific LHCb configuration [32]. The production of mesons is simulated using the dedicated generator BcVegPy [33]. Decays of unstable particles are described by EvtGen [34], in which final-state radiation is generated using Photos [35]. The interaction of the generated particles with the detector, and its response, are implemented using the Geant4 toolkit [36, *Agostinelli:2002hh] as described in Ref. [38]. The underlying interaction is reused multiple times, with an independently generated signal decay for each [39].
@@ 3 Candidate selection and background sources
The optimal requirements for the nonresonant search and for the measurement are referred to subsequently as the and the working points (WPs), respectively. At the WP, the classifier has a combinatorial background rejection power of , whilst retaining of signal decays. At the WP, the classifier has a combinatorial background rejection power of , whilst retaining of signal decays. In both cases, the particle identification requirements have an efficiency around , with a kaon misidentification rate around . To measure branching fraction ratios, the same selection is used for the signal and for the normalisation modes, such that potential systematic biases on the measurement are reduced. For both WPs, each selected event contains only one retained candidate.
@@ 5 Efficiencies and systematic uncertainties
The nonresonant signal decays are simulated with a phase-space distribution of the final-state particles. However, the true distribution of decays is unknown, and there is no clear theoretical guidance for its expectation. A systematic uncertainty is assigned due to the variation of efficiency over , corresponding to the spread of the efficiency in each interval, shown in Fig. 4. The approach of determining the differential branching fraction within bins helps to reduce this uncertainty.
@@ 6 Results
Upper limits on the branching fraction ratios are obtained following the Feldmann–Cousins prescription [55]: pseudoexperiments are generated for various values of and the resulting distribution of measured is used to form confidence belts. Figure 5 shows confidence belts at and confidence level (CL). Table 6 gives the results for and the obtained limits. To assess the impact of the systematic uncertainties, the fits are repeated fixing the nuisance parameters to their central values. Figure 6 summarises the obtained limits on the normalised differential branching fraction.
@@ Acknowledgements
F. Abudinén
T. Ackernley
B. Adeva
M. Adinolfi
P. Adlarson
C. Agapopoulou
C.A. Aidala
Z. Ajaltouni,
S. Akar
K. Akiba
P. Albicocco
J. Albrecht
F. Alessio
M. Alexander
A. Alfonso Albero
Z. Aliouche
P. Alvarez Cartelle
R. Amalric
S. Amato
J.L. Amey
Y. Amhis
L. An
L. Anderlini
M. Andersson
A. Andreianov
P. Andreola
M. Andreotti
D. Andreou
A. Anelli
D. Ao
F. Archilli
M. Argenton
S. Arguedas Cuendis
A. Artamonov
M. Artuso
E. Aslanides
M. Atzeni
B. Audurier
D. Bacher
I. Bachiller Perea
S. Bachmann
M. Bachmayer
J.J. Back
P. Baladron Rodriguez
V. Balagura
W. Baldini
J. Baptista de Souza Leite
M. Barbetti
I. R. Barbosa
R.J. Barlow
S. Barsuk
W. Barter
M. Bartolini
J. Bartz
F. Baryshnikov
J.M. Basels
G. Bassi
B. Batsukh
A. Battig
A. Bay
A. Beck
M. Becker
F. Bedeschi
I.B. Bediaga
A. Beiter,
S. Belin
V. Bellee
K. Belous
I. Belov
I. Belyaev
G. Benane
G. Bencivenni
E. Ben-Haim
A. Berezhnoy
R. Bernet
S. Bernet Andres
C. Bertella
A. Bertolin
C. Betancourt
F. Betti
J. Bex
Ia. Bezshyiko
J. Bhom
M.S. Bieker
N.V. Biesuz
P. Billoir
A. Biolchini
M. Birch
F.C.R. Bishop
A. Bitadze
@@ A. Bizzeti
M.P. Blago
T. Blake
F. Blanc
J.E. Blank
S. Blusk
D. Bobulska
V. Bocharnikov
J.A. Boelhauve
O. Boente Garcia
T. Boettcher
A. Bohare
A. Boldyrev
C.S. Bolognani
R. Bolzonella
N. Bondar
F. Borgato
S. Borghi
M. Borsato
J.T. Borsuk
S.A. Bouchiba
T.J.V. Bowcock
A. Boyer
C. Bozzi
M.J. Bradley,
A. Brea Rodriguez
N. Breer
J. Brodzicka
A. Brossa Gonzalo
J. Brown
D. Brundu
A. Buonaura
L. Buonincontri
A.T. Burke
C. Burr
A. Bursche,
A. Butkevich
J.S. Butter
J. Buytaert
W. Byczynski
S. Cadeddu
H. Cai,
R. Calabrese
L. Calefice
S. Cali
M. Calvi
M. Calvo Gomez
J. Cambon Bouzas
P. Campana
D.H. Campora Perez
A.F. Campoverde Quezada
S. Capelli
L. Capriotti
R. Caravaca-Mora
A. Carbone
L. Carcedo Salgado
R. Cardinale
A. Cardini
P. Carniti
L. Carus,
A. Casais Vidal
R. Caspary
G. Casse
J. Castro Godinez
M. Cattaneo
G. Cavallero
V. Cavallini
S. Celani
J. Cerasoli
D. Cervenkov
S. Cesare
A.J. Chadwick
I. Chahrour
M. Charles
Ph. Charpentier
C.A. Chavez Barajas
M. Chefdeville
C. Chen
S. Chen
Z. Chen
A. Chernov
S. Chernyshenko
V. Chobanova
S. Cholak
M. Chrzaszcz
A. Chubykin
V. Chulikov
P. Ciambrone
M.F. Cicala
X. Cid Vidal
G. Ciezarek
P. Cifra
P.E.L. Clarke
M. Clemencic
H.V. Cliff
J. Closier
J.L. Cobbledick
C. Cocha Toapaxi
V. Coco
J. Cogan
E. Cogneras
L. Cojocariu
P. Collins
T. Colombo
A. Comerma-Montells
L. Congedo
A. Contu
N. Cooke
I. Corredoira
A. Correia
G. Corti
J.J. Cottee Meldrum,
B. Couturier
D.C. Craik
M. Cruz Torres
E. Curras Rivera
R. Currie
C.L. Da Silva
S. Dadabaev
L. Dai
X. Dai
E. Dall’Occo
J. Dalseno
C. D’Ambrosio
J. Daniel
A. Danilina
P. d’Argent
A. Davidson
J.E. Davies
A. Davis
O. De Aguiar Francisco
C. De Angelis
J. de Boer
K. De Bruyn
S. De Capua
M. De Cian
U. De Freitas Carneiro Da Graca
E. De Lucia
J.M. De Miranda
L. De Paula
M. De Serio
D. De Simone
P. De Simone
F. De Vellis
J.A. de Vries
F. Debernardis
D. Decamp
V. Dedu
L. Del Buono
B. Delaney
H.-P. Dembinski
J. Deng
V. Denysenko
O. Deschamps
F. Dettori
B. Dey
P. Di Nezza
I. Diachkov
S. Didenko
S. Ding
V. Dobishuk
A. D. Docheva
A. Dolmatov,
C. Dong
A.M. Donohoe
F. Dordei
A.C. dos Reis
L. Douglas,
A.G. Downes
W. Duan
P. Duda
M.W. Dudek
L. Dufour
V. Duk
P. Durante
M. M. Duras
J.M. Durham
A. Dziurda
A. Dzyuba
S. Easo
E. Eckstein,
U. Egede
A. Egorychev
V. Egorychev
C. Eirea Orro,
S. Eisenhardt
E. Ejopu
S. Ek-In
L. Eklund
M. Elashri
J. Ellbracht
S. Ely
A. Ene
E. Epple
S. Escher
J. Eschle
S. Esen
T. Evans
F. Fabiano
L.N. Falcao
Y. Fan
B. Fang
L. Fantini
M. Faria
K. Farmer
D. Fazzini
L. Felkowski
M. Feng
M. Feo
M. Fernandez Gomez
A.D. Fernez
F. Ferrari
F. Ferreira Rodrigues
S. Ferreres Sole
M. Ferrillo
M. Ferro-Luzzi
S. Filippov
R.A. Fini
M. Fiorini
K.M. Fischer
D.S. Fitzgerald
C. Fitzpatrick
F. Fleuret
M. Fontana
L. F. Foreman
R. Forty
D. Foulds-Holt
M. Franco Sevilla
M. Frank
E. Franzoso
G. Frau
C. Frei
D.A. Friday
L. Frontini
J. Fu
Q. Fuehring
Y. Fujii
T. Fulghesu
E. Gabriel
G. Galati
M.D. Galati
A. Gallas Torreira
D. Galli
S. Gambetta
M. Gandelman
P. Gandini
H. Gao
R. Gao
Y. Gao
Y. Gao
Y. Gao,
M. Garau
L.M. Garcia Martin
P. Garcia Moreno
J. García Pardiñas
B. Garcia Plana,
K. G. Garg
L. Garrido
C. Gaspar
R.E. Geertsema
L.L. Gerken
E. Gersabeck
M. Gersabeck
T. Gershon
Z. Ghorbanimoghaddam,
L. Giambastiani
F. I. Giasemis
V. Gibson
H.K. Giemza
A.L. Gilman
M. Giovannetti
A. Gioventù
P. Gironella Gironell
C. Giugliano
M.A. Giza
E.L. Gkougkousis
F.C. Glaser
V.V. Gligorov
C. Göbel
E. Golobardes
D. Golubkov
A. Golutvin
A. Gomes
S. Gomez Fernandez
F. Goncalves Abrantes
M. Goncerz
G. Gong
J. A. Gooding
I.V. Gorelov
C. Gotti
J.P. Grabowski
L.A. Granado Cardoso
E. Graugés
E. Graverini
L. Grazette
@@ G. Graziani
A. T. Grecu
L.M. Greeven
N.A. Grieser
L. Grillo
S. Gromov
C. Gu
M. Guarise
M. Guittiere
V. Guliaeva
P. A. Günther
A.-K. Guseinov
E. Gushchin
Y. Guz
T. Gys
K. Habermann
T. Hadavizadeh
C. Hadjivasiliou
G. Haefeli
C. Haen
J. Haimberger
M. Hajheidari,
M.M. Halvorsen
P.M. Hamilton
J. Hammerich
Q. Han
X. Han
S. Hansmann-Menzemer
L. Hao
N. Harnew
T. Harrison
M. Hartmann
J. He
K. Heijhoff
F. Hemmer
C. Henderson
R.D.L. Henderson
A.M. Hennequin
K. Hennessy
L. Henry
J. Herd
P. Herrero Gascon
J. Heuel
A. Hicheur
G. Hijano Mendizabal,
D. Hill
S.E. Hollitt
J. Horswill
R. Hou
Y. Hou
N. Howarth,
J. Hu,
J. Hu
W. Hu
X. Hu
W. Huang
W. Hulsbergen
R.J. Hunter
M. Hushchyn
D. Hutchcroft
D. Ilin
P. Ilten
A. Inglessi
A. Iniukhin
A. Ishteev
K. Ivshin
R. Jacobsson
H. Jage
S.J. Jaimes Elles
S. Jakobsen
E. Jans
B.K. Jashal
A. Jawahery
V. Jevtic
E. Jiang
X. Jiang
Y. Jiang
Y. J. Jiang
M. John
D. Johnson
C.R. Jones
T.P. Jones
S. Joshi
B. Jost
N. Jurik
I. Juszczak
D. Kaminaris
S. Kandybei
Y. Kang
M. Karacson
D. Karpenkov
M. Karpov
A. M. Kauniskangas
J.W. Kautz
F. Keizer
D.M. Keller
M. Kenzie
T. Ketel
B. Khanji
A. Kharisova
S. Kholodenko
G. Khreich
T. Kirn
V.S. Kirsebom
O. Kitouni
S. Klaver
N. Kleijne
K. Klimaszewski
M.R. Kmiec
S. Koliiev
L. Kolk
A. Konoplyannikov
P. Kopciewicz
P. Koppenburg
M. Korolev
I. Kostiuk
O. Kot,

### v2

@@
Published in Eur. Phys. J. C 84, 468 (2024)
© 2024 CERN for the benefit of the LHCb collaboration. CC BY 4.0 licence.
@@ 1 Introduction
where indicates a yield, indicates the efficiency determined from simulation with data-driven corrections, and is the known branching fraction of the decay [21]. Here, indicates the differential branching fraction integrated over the range relevant for the given bin.
@@ 2 Detector and simulation
Simulation is used to optimise the event selection procedure, to model the shape of the candidate invariant-mass distributions and to estimate efficiencies accounting for the effects of the detector acceptance, reconstruction and selection criteria. In the simulation, collisions are generated using Pythia [31] with a specific LHCb configuration [32]. The production of mesons is simulated using the dedicated generator BcVegPy [33]. Decays of unstable particles are described by EvtGen [34], in which final-state radiation is generated using Photos [35]. The interaction of the generated particles with the detector, and its response, are implemented using the Geant4 toolkit [36] as described in Ref. [38]. The underlying interaction is reused multiple times, with an independently generated signal decay for each [39].
@@ 3 Candidate selection and background sources
The optimal requirements for the nonresonant search and for the measurement are referred to subsequently as the and the working points (WPs), respectively. At the WP, the classifier has a combinatorial background rejection power of , whilst retaining of signal decays. At the WP, the classifier has a combinatorial background rejection power of , whilst retaining of signal decays. In both cases, the particle identification requirements have a pion efficiency around , with a kaon misidentification rate around . The particle identification requirements applied on the muon candidates have an efficiency around , with a pion misidentification rate below . To measure branching fraction ratios, the same selection is used for the signal and for the normalisation modes, such that potential systematic biases on the measurement are reduced. For both WPs, each selected event contains only one retained candidate.
@@ 5 Efficiencies and systematic uncertainties
The nonresonant signal decays are simulated with a phase-space distribution of the final-state particles. However, the true distribution of decays is unknown, and there is no clear theoretical guidance for its expectation. A systematic uncertainty is assigned due to the change of efficiency over , shown in Fig. 4. The uncertainty is taken to be the root-mean-square variation of the distribution of the efficiency values in narrow bins within each interval. The approach of determining the differential branching fraction within bins helps to reduce this uncertainty.
@@ 6 Results
Upper limits on the branching fraction ratios are obtained following the Feldmann–Cousins prescription [55]: pseudoexperiments are generated for various values of and the resulting distribution of measured is used to form confidence belts. Figure 5 shows confidence belts at and confidence level (CL). Table 6 gives the results for and the obtained limits. To assess the impact of the systematic uncertainties, the fits are repeated fixing the nuisance parameters to their central values. Figure 6 summarises the obtained limits on the normalised differential branching fraction. As further checks the procedure is repeated restricting the signal yield to positive values, or performing nonextended maximum-likelihood fits. No significant changes in the obtained upper limits are found.
@@ Acknowledgements
F. Abudinén54
T. Ackernley58
B. Adeva44
M. Adinolfi52
P. Adlarson78
C. Agapopoulou46
C.A. Aidala79
Z. Ajaltouni11,
S. Akar63
K. Akiba35
P. Albicocco25
J. Albrecht17
F. Alessio46
M. Alexander57
A. Alfonso Albero43
Z. Aliouche60
P. Alvarez Cartelle53
R. Amalric15
S. Amato3
J.L. Amey52
Y. Amhis13,46
L. An6
L. Anderlini24
M. Andersson48
A. Andreianov41
P. Andreola48
M. Andreotti23
D. Andreou66
A. Anelli28,o
D. Ao7
F. Archilli34,u
M. Argenton23
S. Arguedas Cuendis9
A. Artamonov41
M. Artuso66
E. Aslanides12
M. Atzeni62
B. Audurier14
D. Bacher61
I. Bachiller Perea10
S. Bachmann19
M. Bachmayer47
J.J. Back54
P. Baladron Rodriguez44
V. Balagura14
W. Baldini23
J. Baptista de Souza Leite2
M. Barbetti24,l
I. R. Barbosa67
R.J. Barlow60
S. Barsuk13
W. Barter56
M. Bartolini53
J. Bartz66
F. Baryshnikov41
J.M. Basels16
G. Bassi32,r
B. Batsukh5
A. Battig17
A. Bay47
A. Beck54
M. Becker17
F. Bedeschi32
I.B. Bediaga2
A. Beiter66,
S. Belin44
V. Bellee48
K. Belous41
I. Belov26
I. Belyaev41
G. Benane12
G. Bencivenni25
E. Ben-Haim15
A. Berezhnoy41
R. Bernet48
S. Bernet Andres42
C. Bertella60
A. Bertolin30
C. Betancourt48
F. Betti56
J. Bex53
Ia. Bezshyiko48
J. Bhom38
M.S. Bieker17
N.V. Biesuz23
P. Billoir15
A. Biolchini35
M. Birch59
F.C.R. Bishop10
A. Bitadze60
@@ A. Bizzeti
M.P. Blago53
T. Blake54
F. Blanc47
J.E. Blank17
S. Blusk66
D. Bobulska57
V. Bocharnikov41
J.A. Boelhauve17
O. Boente Garcia14
T. Boettcher63
A. Bohare56
A. Boldyrev41
C.S. Bolognani76
R. Bolzonella23,k
N. Bondar41
F. Borgato30,46
S. Borghi60
M. Borsato28,o
J.T. Borsuk38
S.A. Bouchiba47
T.J.V. Bowcock58
A. Boyer46
C. Bozzi23
M.J. Bradley59,
A. Brea Rodriguez44
N. Breer17
J. Brodzicka38
A. Brossa Gonzalo44
J. Brown58
D. Brundu29
A. Buonaura48
L. Buonincontri30
A.T. Burke60
C. Burr46
A. Bursche69,
A. Butkevich41
J.S. Butter53
J. Buytaert46
W. Byczynski46
S. Cadeddu29
H. Cai71,
R. Calabrese23,k
L. Calefice17
S. Cali25
M. Calvi28,o
M. Calvo Gomez42
J. Cambon Bouzas44
P. Campana25
D.H. Campora Perez76
A.F. Campoverde Quezada7
S. Capelli28,o
L. Capriotti23
R. Caravaca-Mora9
A. Carbone22,i
L. Carcedo Salgado44
R. Cardinale26,m
A. Cardini29
P. Carniti28,o
L. Carus19,
A. Casais Vidal62
R. Caspary19
G. Casse58
J. Castro Godinez9
M. Cattaneo46
G. Cavallero23
V. Cavallini23,k
S. Celani19
J. Cerasoli12
D. Cervenkov61
S. Cesare27,n
A.J. Chadwick58
I. Chahrour79
M. Charles15
Ph. Charpentier46
C.A. Chavez Barajas58
M. Chefdeville10
C. Chen12
S. Chen5
Z. Chen7
A. Chernov38
S. Chernyshenko50
V. Chobanova44,y
S. Cholak47
M. Chrzaszcz38
A. Chubykin41
V. Chulikov41
P. Ciambrone25
M.F. Cicala54
X. Cid Vidal44
G. Ciezarek46
P. Cifra46
P.E.L. Clarke56
M. Clemencic46
H.V. Cliff53
J. Closier46
J.L. Cobbledick60
C. Cocha Toapaxi19
V. Coco46
J. Cogan12
E. Cogneras11
L. Cojocariu40
P. Collins46
T. Colombo46
A. Comerma-Montells43
L. Congedo21
A. Contu29
N. Cooke57
I. Corredoira 44
A. Correia15
G. Corti46
J.J. Cottee Meldrum52,
B. Couturier46
D.C. Craik48
M. Cruz Torres2,g
E. Curras Rivera47
R. Currie56
C.L. Da Silva65
S. Dadabaev41
L. Dai68
X. Dai6
E. Dall’Occo17
J. Dalseno44
C. D’Ambrosio46
J. Daniel11
A. Danilina41
P. d’Argent21
A. Davidson54
J.E. Davies60
A. Davis60
O. De Aguiar Francisco60
C. De Angelis29,j
J. de Boer35
K. De Bruyn75
S. De Capua60
M. De Cian19,46
U. De Freitas Carneiro Da Graca2,b
E. De Lucia25
J.M. De Miranda2
L. De Paula3
M. De Serio21,h
D. De Simone48
P. De Simone25
F. De Vellis17
J.A. de Vries76
F. Debernardis21,h
D. Decamp10
V. Dedu12
L. Del Buono15
B. Delaney62
H.-P. Dembinski17
J. Deng8
V. Denysenko48
O. Deschamps11
F. Dettori29,j
B. Dey74
P. Di Nezza25
I. Diachkov41
S. Didenko41
S. Ding66
V. Dobishuk50
A. D. Docheva57
A. Dolmatov41,
C. Dong4
A.M. Donohoe20
F. Dordei29
A.C. dos Reis2
L. Douglas57,
A.G. Downes10
W. Duan69
P. Duda77
M.W. Dudek38
L. Dufour46
V. Duk31
P. Durante46
M. M. Duras77
J.M. Durham65
A. Dziurda38
A. Dzyuba41
S. Easo55,46
E. Eckstein73,
U. Egede1
A. Egorychev41
V. Egorychev41
C. Eirea Orro44,
S. Eisenhardt56
E. Ejopu60
S. Ek-In47
L. Eklund78
M. Elashri63
J. Ellbracht17
S. Ely59
A. Ene40
E. Epple63
S. Escher16
J. Eschle48
S. Esen19
T. Evans60
F. Fabiano29,j,46
L.N. Falcao2
Y. Fan7
B. Fang71,13
L. Fantini31,q
M. Faria47
K. Farmer56
D. Fazzini28,o
L. Felkowski77
M. Feng5,7
M. Feo46
M. Fernandez Gomez44
A.D. Fernez64
F. Ferrari22
F. Ferreira Rodrigues3
S. Ferreres Sole35
M. Ferrillo48
M. Ferro-Luzzi46
S. Filippov41
R.A. Fini21
M. Fiorini23,k
K.M. Fischer61
D.S. Fitzgerald79
C. Fitzpatrick60
F. Fleuret14
M. Fontana22
L. F. Foreman60
R. Forty46
D. Foulds-Holt53
M. Franco Sevilla64
M. Frank46
E. Franzoso23,k
G. Frau19
C. Frei46
D.A. Friday60
L. Frontini27
J. Fu7
Q. Fuehring17
Y. Fujii1
T. Fulghesu15
E. Gabriel35
G. Galati21,h
M.D. Galati35
A. Gallas Torreira44
D. Galli22,i
S. Gambetta56
M. Gandelman3
P. Gandini27
H. Gao7
R. Gao61
Y. Gao8
Y. Gao6
Y. Gao8,
M. Garau29,j
L.M. Garcia Martin47
P. Garcia Moreno43
J. García Pardiñas46
B. Garcia Plana44,
K. G. Garg8
L. Garrido43
C. Gaspar46
R.E. Geertsema35
L.L. Gerken17
E. Gersabeck60
M. Gersabeck60
T. Gershon54
Z. Ghorbanimoghaddam52,
L. Giambastiani30
F. I. Giasemis15,e
V. Gibson53
H.K. Giemza39
A.L. Gilman61
M. Giovannetti25
A. Gioventù43
P. Gironella Gironell43
C. Giugliano23,k
M.A. Giza38
E.L. Gkougkousis59
F.C. Glaser13,19
V.V. Gligorov15
C. Göbel67
E. Golobardes42
D. Golubkov41
A. Golutvin59,41,46
A. Gomes2,a,†
S. Gomez Fernandez43
F. Goncalves Abrantes61
M. Goncerz38
G. Gong4
J. A. Gooding17
I.V. Gorelov41
C. Gotti28
J.P. Grabowski73
L.A. Granado Cardoso46
E. Graugés43
E. Graverini47,s
L. Grazette54
@@ G. Graziani
A. T. Grecu40
L.M. Greeven35
N.A. Grieser63
L. Grillo57
S. Gromov41
C. Gu14
M. Guarise23
M. Guittiere13
V. Guliaeva41
P. A. Günther19
A.-K. Guseinov41
E. Gushchin41
Y. Guz6,41,46
T. Gys46
K. Habermann73
T. Hadavizadeh1
C. Hadjivasiliou64
G. Haefeli47
C. Haen46
J. Haimberger46
M. Hajheidari46,
M.M. Halvorsen46
P.M. Hamilton64
J. Hammerich58
Q. Han8
X. Han19
S. Hansmann-Menzemer19
L. Hao7
N. Harnew61
T. Harrison58
M. Hartmann13
J. He7,c
K. Heijhoff35
F. Hemmer46
C. Henderson63
R.D.L. Henderson1,54
A.M. Hennequin46
K. Hennessy58
L. Henry47
J. Herd59
P. Herrero Gascon19
J. Heuel16
A. Hicheur3
G. Hijano Mendizabal48,
D. Hill47
S.E. Hollitt17
J. Horswill60
R. Hou8
Y. Hou10
N. Howarth58,
J. Hu19,
J. Hu69
W. Hu6
X. Hu4
W. Huang7
W. Hulsbergen35
R.J. Hunter54
M. Hushchyn41
D. Hutchcroft58
D. Ilin41
P. Ilten63
A. Inglessi41
A. Iniukhin41
A. Ishteev41
K. Ivshin41
R. Jacobsson46
H. Jage16
S.J. Jaimes Elles45,72
S. Jakobsen46
E. Jans35
B.K. Jashal45
A. Jawahery64,46
V. Jevtic17
E. Jiang64
X. Jiang5,7
Y. Jiang7
Y. J. Jiang6
M. John61
D. Johnson51
C.R. Jones53
T.P. Jones54
S. Joshi39
B. Jost46
N. Jurik46
I. Juszczak38
D. Kaminaris47
S. Kandybei49
Y. Kang4
M. Karacson46
D. Karpenkov41
M. Karpov41
A. M. Kauniskangas47
J.W. Kautz63
F. Keizer46
D.M. Keller66
M. Kenzie53
T. Ketel35
B. Khanji66
A. Kharisova41
S. Kholodenko32
G. Khreich13
T. Kirn16
V.S. Kirsebom47
O. Kitouni62
S. Klaver36
N. Kleijne32,r
K. Klimaszewski39
M.R. Kmiec39
S. Koliiev50
L. Kolk17
A. Konoplyannikov41
P. Kopciewicz37,46
P. Koppenburg35
M. Korolev41
I. Kostiuk35
O. Kot50,